if a rectangle table is 2 times as long as it is wide and the area is 108ft^2, what is the length and width?
Answer by amoresroy(361)

Let x = with of rectangle
2x = lenght of rectangle
A = x (2x)
x (2x) = 108
2x^2 = 108
x^2 = 54

x = suare root of 54
x = 7.35 ft
2x = 14.7 ft